### v1

@@ Abstract
Keywords:
Natural Language Generation Evaluation Metrics Large Language Models Text Quality Assessment
@@ 1 Introduction
Evaluating the quality of text generated by large language models (LLMs) remains an open problem within the field of natural language generation (NLG). Traditional evaluation metrics, such as BLEU [6], ROUGE [4], and METEOR [1], often show limited correlation with human judgments, especially in tasks that require creativity and diversity, such as dialogue generation and summarization. Despite advancements in LLMs, which can produce high-quality, fluent texts that closely mirror human writing, the challenge lies in accurately assessing these outputs quality.
Recent approaches utilizing LLMs as text quality evaluators have shown promise, yet they still fall short in achieving reliable alignment with human judgments [5, 3]. This ongoing challenge underscores the necessity for more effective and scalable evaluation frameworks that can bridge the gap between automated metrics and human evaluators, ensuring that the outputs of NLG systems meet the desired standards of coherence, relevance, and overall quality [5].
@@
We evaluate Check-Eval in two main scenarios, both based on human judgments: (1) the Portuguese Legal Semantic Textual Similarity dataset [7], a benchmark dataset for evaluating the semantic similarity of legal texts in Portuguese, and (2) the SummEval dataset [2], a benchmark dataset for text summarization evaluation. Our experiments demonstrate that Check-Eval achieves higher correlations with human judgments compared to existing metrics, such as G-Eval and GPTScore, highlighting its potential as a more reliable and effective evaluation framework for NLG tasks. Additionally, we show that Check-Eval can identify specific areas of improvement in the generated summaries, providing valuable feedback for model development and refinement.
@@ 2 Related Work
The evaluation of automatically generated text has been a persistent challenge in the NLG field. Traditional metrics such as BLEU, ROUGE, and METEOR have been extensively used but have shown limitations in aligning with human judgment, particularly in tasks requiring creativity and nuance [2]. In recent years, more sophisticated evaluation frameworks leveraging LLMs have been proposed to address these shortcomings. Two recent methods are GPTScore [3] and G-Eval [5].
GPTScore is a framework that utilizes generative pre-trained transformers (GPTs) and other language models to evaluate NLG outputs without relying on reference texts. The core idea is to assess the probability that the LLM assigns to the generated text, under the assumption that higher probabilities indicate higher quality. Fu et al. (2023) [3] demonstrated that GPTScore could achieve better correlations with human judgments compared to traditional metrics, especially in open-ended tasks such as dialogue generation and creative writing. However, despite its advancements, the method lacks interpretability and may be biased towards texts similar to those seen during the model’s training phase.
G-Eval is another recent approach that leverages the capabilities of LLMs, specifically GPT-4, to improve NLG evaluation. Proposed by Liu et al. (2023) [5], G-Eval introduces a chain-of-thought (CoT) [8] paradigm where the evaluation process is guided by detailed intermediate steps generated by the LLM. This method has shown improvements in correlation with human evaluations, particularly in tasks such as text summarization and dialogue generation. To address issues with score distribution and variability, G-Eval employs a self-consistency strategy. Specifically, it generates multiple samples ($n=20$) using different decoding parameters and averages the evaluation scores across these samples. This approach helps mitigate two key problems: the dominance of a single score (such as 3 on a 1-5 scale) and the tendency of LLMs to output integer scores, even when decimal values are requested. By using self-consistency, G-Eval captures more subtle differences between generated texts and provides a more reliable assessment of text quality.
@@ 3.1.1 Reference-Guided and Candidate-Guided Checklist Generation
Relevance definition: selection of important content from the source. The summary should include only important information from the source document. Annotators were instructed to penalize summaries that contained redundancies and excess information.
You must follow the following rules:
1. The checklist should contain yes/no questions.
2. Checklist items must be self-contained.
3. Focus on the main concepts and avoid including minor details.
4. Each item should represent a unique concept or element from the text.
5. Avoid repetition and overlapping of concepts in the checklist items.
6. The checklist should be comprehensive but not exhaustive, aiming for clarity and brevity.
7. Generate as many items as you think are necessary
[User]
{source text}
@@ 7. Generate as many items as you think are necessary
Source Document Excerpt: "Climate change refers to long-term shifts and alterations in temperature and weather patterns. These changes may be natural, such as through variations in the solar cycle. However, since the 1800s, human activities have been the main driver of climate change, primarily due to the burning of fossil fuels like coal, oil, and gas."
Generated Checklist: 1. Does the summary mention that climate change refers to long-term shifts in temperature and weather patterns? 2. Does the summary state that these changes can be natural, such as variations in the solar cycle? 3. Does the summary highlight that human activities have been the main driver of climate change since the 1800s? 4. Does the summary specify that burning fossil fuels like coal, oil, and gas is a primary cause?
@@ 3.2 Checklist Evaluation
Consistency definition: the factual alignment between the summary and the summarized source. A factually consistent summary contains only statements that are entailed by the source document. Annotators were also asked to penalize summaries that contained hallucinated facts.
Checklist:
{generated checklist}
[User]
{candidate summary}
@@ 4.1 Portuguese Legal Semantic Textual Similarity
We evaluated Check-Eval using the Portuguese Legal Semantic Textual Similarity [7] dataset, a benchmark dataset for evaluating the semantic similarity of legal texts in Portuguese. The dataset consists of pairs of legal documents, each annotated with a similarity score. The dataset provides two versions of the annotations: one performed by legal experts on 32 pairs of legal documents and another automated annotation using heuristics based on legal cases metadata. The overall correlation between the human and automated annotations is 0.45 (Pearson correlation coefficient) and 0.43 (Spearman correlation coefficient). We experimented with the 32 pairs annotated by legal experts and a subset of 100 pairs annotated by the automated method randomly selected from the dataset.
@@ 4.2 SummEval Dataset
We also evaluated Check-Eval in a reference-free evaluation scenario using the SummEval dataset [2]. The SummEval dataset consists of automatically generated summaries for news articles from the CNN/DailyMail dataset, along with human annotations for the quality dimensions of coherence, consistency, fluency, and relevance. This dataset provides a comprehensive benchmark for comparing the performance of different evaluation metrics against human judgments. In this case, we evaluated the Criterion-Guided variation of Check-Eval using the SummEval dataset.
@@ 4.2.1 Experimental Settings
For our experiments, we used the GPT-4 model to perform both checklist generation and checklist evaluation. We prompted GPT-4 to generate a checklist of key points based on the evaluation criteria provided in the SummEval paper [2]. The following steps outline the experimental setup:
@@
    Checklist Generation: For each of the evaluation criteria (consistency, relevance, coherence, and fluency), we prompted GPT-4 to generate a checklist of key points based on the criteria definitions provided in the SummEval paper [2].
@@ 5.1 Portuguese Legal Semantic Textual Similarity
We also evaluated Check-Eval using the automated annotations from the dataset. Table 2 presents the results for a sub-sample of 100 pairs automatically annotated by the heuristics proposed in [7]. The results show that Check-Eval achieves higher correlations with the automated annotations compared to the human annotations. For example, the reference-guided variation shows a Pearson correlation of 0.53 for consistency with automated annotations, whereas it shows 0.45 with human annotations. This suggests that the automated annotations align more closely with the systematic evaluation approach of Check-Eval.
Table 2: Check-Eval results for Portuguese Legal Semantic Textual Similarity. The table presents ($\rho$) Pearson and ($\rho_{s}$) Spearman correlations computed for a sub-sample of 100 pairs automatically annotated by the heuristics proposed in [7].

### v2

@@ 1 Introduction
Evaluating the quality of text generated by large language models (LLMs) remains an open problem within the field of natural language generation (NLG). Traditional evaluation metrics, such as BLEU [Papineni et al. 2002], ROUGE [Lin 2004], and METEOR [Banerjee and Lavie 2005], often show limited correlation with human judgments, especially in tasks that require creativity and diversity, such as dialogue generation and summarization. Despite advancements in LLMs, which can produce high-quality, fluent texts that closely mirror human writing, the challenge lies in accurately assessing these outputs quality.
Recent approaches utilizing LLMs as text quality evaluators have shown promise, yet they still fall short in achieving reliable alignment with human judgments [Liu et al. 2023, Fu et al. 2023]. This ongoing challenge underscores the necessity for more effective and scalable evaluation frameworks that can bridge the gap between automated metrics and human evaluators, ensuring that the outputs of NLG systems meet the desired standards of coherence, relevance, and overall quality [Liu et al. 2023].
@@
We evaluate Check-Eval in two main scenarios, both based on human judgments: (1) the Portuguese Legal Semantic Textual Similarity dataset [da Silva Junior et al. 2024], a benchmark dataset for evaluating the semantic similarity of legal texts in Portuguese, and (2) the SummEval dataset [Fabbri et al. 2021], a benchmark dataset for text summarization evaluation. Our experiments demonstrate that Check-Eval achieves higher correlations with human judgments compared to existing metrics, such as G-Eval and GPTScore, highlighting its potential as a more reliable and effective evaluation framework for NLG tasks. Additionally, we show that Check-Eval can identify specific areas of improvement in the generated summaries, providing valuable feedback for model development and refinement.
@@ 2 Related Work
The evaluation of automatically generated text has been a persistent challenge in the NLG field. Traditional metrics such as BLEU, ROUGE, and METEOR have been extensively used but have shown limitations in aligning with human judgment, particularly in tasks requiring creativity and nuance [Fabbri et al. 2021]. In recent years, more sophisticated evaluation frameworks leveraging LLMs have been proposed to address these shortcomings. Two recent methods are GPTScore [Fu et al. 2023] and G-Eval [Liu et al. 2023].
GPTScore is a framework that utilizes generative pre-trained transformers (GPTs) and other language models to evaluate NLG outputs without relying on reference texts. The core idea is to assess the probability that the LLM assigns to the generated text, under the assumption that higher probabilities indicate higher quality. Fu et al. (2023) [Fu et al. 2023] demonstrated that GPTScore could achieve better correlations with human judgments compared to traditional metrics, especially in open-ended tasks such as dialogue generation and creative writing. However, despite its advancements, the method lacks interpretability and may be biased towards texts similar to those seen during the model’s training phase.
G-Eval is another recent approach that leverages the capabilities of LLMs, specifically GPT-4, to improve NLG evaluation. Proposed by Liu et al. (2023) [Liu et al. 2023], G-Eval introduces a chain-of-thought (CoT) [Wei et al. 2022] paradigm where the evaluation process is guided by detailed intermediate steps generated by the LLM. This method has shown improvements in correlation with human evaluations, particularly in tasks such as text summarization and dialogue generation. To address issues with score distribution and variability, G-Eval employs a self-consistency strategy. Specifically, it generates multiple samples ($n=20$) using different decoding parameters and averages the evaluation scores across these samples. This approach helps mitigate two key problems: the dominance of a single score (such as 3 on a 1-5 scale) and the tendency of LLMs to output integer scores, even when decimal values are requested. By using self-consistency, G-Eval captures more subtle differences between generated texts and provides a more reliable assessment of text quality.
@@ 3.1.1 Reference-Guided and Candidate-Guided Checklist Generation
Relevance definition: selection of important content from the source. The summary should include only important information from the source document. Annotators were instructed to penalize summaries that contained redundancies and excess information. You must follow the following rules: 1. The checklist should contain yes/no questions. 2. Checklist items must be self-contained. 3. Focus on the main concepts and avoid including minor details. 4. Each item should represent a unique concept or element from the text. 5. Avoid repetition and overlapping of concepts in the checklist items. 6. The checklist should be comprehensive but not exhaustive, aiming for clarity and brevity. 7. Generate as many items as you think are necessary [User] {source text}
@@
Source Document Excerpt: "Climate change refers to long-term shifts and alterations in temperature and weather patterns. These changes may be natural, such as through variations in the solar cycle. However, since the 1800s, human activities have been the main driver of climate change, primarily due to the burning of fossil fuels like coal, oil, and gas." Generated Checklist: 1. Does the summary mention that climate change refers to long-term shifts in temperature and weather patterns? 2. Does the summary state that these changes can be natural, such as variations in the solar cycle? 3. Does the summary highlight that human activities have been the main driver of climate change since the 1800s? 4. Does the summary specify that burning fossil fuels like coal, oil, and gas is a primary cause?
@@ 3.2 Checklist Evaluation
Consistency definition: the factual alignment between the summary and the summarized source. A factually consistent summary contains only statements that are entailed by the source document. Annotators were also asked to penalize summaries that contained hallucinated facts. Checklist: {generated checklist} [User] {candidate summary}
@@ 4.1 Portuguese Legal Semantic Textual Similarity
We evaluated Check-Eval using the Portuguese Legal Semantic Textual Similarity [da Silva Junior et al. 2024] dataset, a benchmark dataset for evaluating the semantic similarity of legal texts in Portuguese. The dataset consists of pairs of legal documents, each annotated with a similarity score. The dataset provides two versions of the annotations: one performed by legal experts on 32 pairs of legal documents and another automated annotation using heuristics based on legal cases metadata. The overall correlation between the human and automated annotations is 0.45 (Pearson correlation coefficient) and 0.43 (Spearman correlation coefficient). We experimented with the 32 pairs annotated by legal experts and a subset of 100 pairs annotated by the automated method randomly selected from the dataset.
@@ 4.2 SummEval Dataset
We also evaluated Check-Eval in a reference-free evaluation scenario using the SummEval dataset [Fabbri et al. 2021]. The SummEval dataset consists of automatically generated summaries for news articles from the CNN/DailyMail dataset, along with human annotations for the quality dimensions of coherence, consistency, fluency, and relevance. This dataset provides a comprehensive benchmark for comparing the performance of different evaluation metrics against human judgments. In this case, we evaluated the Criterion-Guided variation of Check-Eval using the SummEval dataset.
@@ 4.2.1 Experimental Settings
For our experiments, we used the GPT-4 model to perform both checklist generation and checklist evaluation. We prompted GPT-4 to generate a checklist of key points based on the evaluation criteria provided in the SummEval paper [Fabbri et al. 2021]. The following steps outline the experimental setup:
@@
    Checklist Generation: For each of the evaluation criteria (consistency, relevance, coherence, and fluency), we prompted GPT-4 to generate a checklist of key points based on the criteria definitions provided in the SummEval paper [Fabbri et al. 2021].
@@ 5.1 Portuguese Legal Semantic Textual Similarity
We also evaluated Check-Eval using the automated annotations from the dataset. Table 2 presents the results for a sub-sample of 100 pairs automatically annotated by the heuristics proposed in [da Silva Junior et al. 2024]. The results show that Check-Eval achieves higher correlations with the automated annotations compared to the human annotations. For example, the reference-guided variation shows a Pearson correlation of 0.53 for consistency with automated annotations, whereas it shows 0.45 with human annotations. This suggests that the automated annotations align more closely with the systematic evaluation approach of Check-Eval.
Table 2: Check-Eval results for Portuguese Legal Semantic Textual Similarity. The table presents ($\rho$) Pearson and ($\rho_{s}$) Spearman correlations computed for a sub-sample of 100 pairs automatically annotated by the heuristics proposed in [da Silva Junior et al. 2024].